OpenStudy (anonymous):

sytem of linear equations solve x+y+4z=-11 x+y+6z=-15 x+8y+8z=-26

hero (hero):

Hint, isolate z in the first equation, then substitute the expression for z into the last two equations.

OpenStudy (anonymous):

We shall begin by getting rid of the x. x+y+4z=-11 -1(x+y+6z=-15) <-- Must make it negative to cancel it out x+y+4z=-11 -x-y-6z=15 <-- New equation after you multiply the -1 = -2z=4 z=-2 Now you solve for the other equation. x+y+4z=-11 -x-8y-8z=26 Answer= -7y-4z=15 <-- Plug in -2 for z -7y+8=15 y=-1 plug in -2 for z and -1 for y in any of the first 3 starting equations to solve for x.